OpenStudy (anonymous):

Solve the equation sin theta = 0.3 to the nearest tenth. Use the restrictions 90degrees < theta < 180degrees.

OpenStudy (saifoo.khan):

Basic angle = 14.46* Sine is positive in first and second quadrant. Hence, Angle 1: Angle 2: 14.46 180-14.46

OpenStudy (anonymous):

A)\[\theta=107.5\] B)\[\theta=162.5\] C)\[\theta=197.5\] D)\[\theta=17.5\]

OpenStudy (saifoo.khan):

it must be b. And it should be 165.5?

OpenStudy (anonymous):

I dont know. The paper says 162.5

OpenStudy (saifoo.khan):

Oh right. I did a typo. Basic angle = 17.46* Sine is positive in first and second quadrant. Hence, Angle 1: Angle 2: 17.46 180-17.46

OpenStudy (saifoo.khan):

And since the range is given, we will only consider angle 2.